OpenStudy (anonymous):

Use a graphing calculator to solve the equation -3 cos t = 1 in the interval from 0 to 2p. Round to the nearest hundredth

OpenStudy (alexwee123):

i dont see why you'd need a graphing calculator for this o_0 a scientific + knowledge of trig is more than enough so -3cost=1 cost=-1/3 \[t=\cos^{-1} (\frac{ -1 }{ 3 })\] t=1.91,1.91+pi

OpenStudy (anonymous):

and then i round it to 2 + pi?

OpenStudy (alexwee123):

nope t=1.91,1.91+pi or t={1.91,5.05} are already the answers

OpenStudy (anonymous):

right thanks!

OpenStudy (alexwee123):

np :)

OpenStudy (anonymous):

hey do you think you can help me with one more question?